# QWHA: Quantization-Aware Walsh-Hadamard Adaptation for Parameter-Efficient Fine-Tuning on Large Language Models

### v3

@@ C.1 Objective Function
We reduce the original optimization objective in Equation 5 to the form in Equation 6, following the approach of Frantar et al. (2023); Deng et al. (2025). Using the notations from Section 3.2, the reduction proceeds as follows:
$\lVert\Delta{\bm{W}}_{Q}{\bm{X}}-{\bm{F}}{\bm{H}}^{-1}{\bm{X}}\rVert_{F}^{2}$ (27)
$=$ $\lVert(\Delta{\bm{W}}_{Q}-{\bm{F}}{\bm{H}}^{-1}){\bm{X}}\rVert_{F}^{2}$ (28)
@@
where ${\bm{R}}={\bm{U}}\mathbf{\Sigma}^{1/2}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{in}}}$ is an invertible square root of the Hessian Gram matrix ${\bm{X}}{\bm{X}}^{\top}$, such that ${\bm{X}}{\bm{X}}^{\top}={\bm{U}}\mathbf{\Sigma}{\bm{U}}^{\top}$. Following Deng et al. (2025), we add a small regularization term $\lambda=0.0001\cdot\operatorname{\mathrm{tr}}({\bm{X}}{\bm{X}}^{\top})/d_{\text{in}}$ to the diagonal if $\bm{R}$} is not originally invertible. This reduction allows us to replace $\bm{X}$} with $\bm{R}$}, enabling efficient and effective calibration using multiple input data points. Rather than solving the optimization problem separately for each sample $\bm{X}$}, we can accumulate the contribution of activations via $\bm{R}$} and solve a single reduced problem.
@@ D.1 Fine-tuning Hyperparameters
We follow the hyperparameter settings adapted from Gao et al. (2024b) and Deng et al. (2025). Training is performed using the AdamW optimizer (Loshchilov & Hutter, 2017). Table 8 reports the key settings, including minibatch size, weight decay, dropout ratio, learning rate scheduler, maximum sequence length, number of training epochs, warmup ratio, and the adapter scaling factor $\alpha$, which can be applied to adapters in Equation 2 through Equation 4. The scaling factors for low-rank and FT-based adapters are defined so that the $\ell_{2}$ norm of the gradients applied to their respective parameters is aligned. The learning rates for each combination of model, task, method, and bit-width are summarized in Table 9. We note that 128 sequences of length 2048, randomly sampled from the WikiText-2 (Merity et al., 2016) dataset, are used as a calibration set for quantization and adapter initialization, as these processes are robust to the choice of dataset (Frantar et al., 2023; Zhang et al., 2024; Deng et al., 2025).
@@ D.2 Commonsense Question-Answering Benchmark
We present the detailed accuracy scores on the commonsense question answering (CSQA) benchmark in this section, while the averaged scores are reported in Section 4. The results for each of the seven individual tasks, are provided in Table 10 and Table D.2, corresponding to Table 4 and Table 4.1, respectively.
@@
D.3 Ablation on Quantization Group Size
@@
Table 12: GSM8k accuracy (%) of QWHA on LLaMA-3.2-3B with 2-bit quantization using various quantization group sizes. We adopt a group size of 64 for the main experiments.
@@
D.4 Training Time of Adapters
@@
D.5 Training Memory Usage of Adapters

### v4

@@ C.1 Objective Function
We reduce the original optimization objective in Equation 5 to the form in Equation 6, following the approach of frantar2023gptq; deng2025cloq. Using the notations from Section 3.2, the reduction proceeds as follows:
$\lVert\Delta{\bm{W}}_{Q}{\bm{X}}-{\bm{F}}{\bm{H}}^{-1}{\bm{X}}\rVert_{F}^{2}=$ $\lVert(\Delta{\bm{W}}_{Q}-{\bm{F}}{\bm{H}}^{-1}){\bm{X}}\rVert_{F}^{2}$ (27)
@@
where ${\bm{R}}={\bm{U}}\mathbf{\Sigma}^{1/2}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{in}}}$ is an invertible square root of the Hessian Gram matrix ${\bm{X}}{\bm{X}}^{\top}$. This term is obtained by applying the SVD ${\bm{X}}{\bm{X}}^{\top}={\bm{U}}\mathbf{\Sigma}{\bm{U}}^{\top}$, where $\mathbf{\Sigma}$ contains the eigenvalues on the diagonal and $\bm{U}$} is the matrix of orthonormal eigenvectors. Following deng2025cloq, we add a small regularization term $\lambda=0.0001\cdot\operatorname{\mathrm{tr}}({\bm{X}}{\bm{X}}^{\top})/d_{\text{in}}$ to the diagonal if $\bm{R}$} is not originally invertible. This reduction allows us to replace $\bm{X}$} with $\bm{R}$}, enabling efficient and effective calibration using multiple input data points. Rather than solving the optimization problem separately for each sample $\bm{X}$}, we can accumulate the contribution of activations via $\bm{R}$} and solve a single reduced problem.
@@ D.1 Fine-tuning Hyperparameters
We follow the hyperparameter settings adapted from gao2024fourierft and deng2025cloq. Training is performed using the AdamW optimizer (loshchilov2017adamw). Table 8 reports the key settings, including minibatch size, weight decay, dropout ratio, learning rate scheduler, maximum sequence length, number of training epochs, warmup ratio, and the adapter scaling factor $\alpha$, which can be applied to adapters in Equation 2 through Equation 4. Due to an implementation detail in our codebase, the explicitly specified scaling factor is internally divided by the layer input dimension $d_{\text{in}}$. As a consequence, the actual scaling applied during training is $\alpha_{\text{effective}}=\alpha_{\text{explicit}}/d_{\text{in}}$. To match the effective scaling used in CLoQ (i.e., $\alpha_{\text{effective}}\approx 1.0$), we set the explicit scaling factor to $\alpha_{\text{explicit}}=4000$, which is close to the typical input dimensions of our models (3072 for LLaMA-3.2-3B and 4096 for LLaMA-3.1-8B and Mistral-7B-v0.3). Under this implementation, the effective scaling becomes $\alpha_{\text{effective}}=4000/d_{\text{in}}\approx 1.0$, ensuring consistent gradient scaling between low-rank. The learning rates for each combination of model, task, method, and bit-width are summarized in Table 9. We note that 128 sequences of length 2048, randomly sampled from the WikiText-2 (wiki2016merity) dataset, are used as a calibration set for quantization and adapter initialization, as these processes are robust to the choice of dataset (frantar2023gptq; zhang2024magrweightmagnitudereduction; deng2025cloq). The total number of parameters for $P(r=64)$ is reported in Table 10, broken down by each projection, the layers containing these projections, and the entire model.
@@
Table 10: Total number of parameters at $P(r=64)$ for each projection, layer, and model.
Model q_proj k_proj v_proj o_proj gate_proj up_proj down_proj per-layer per-model
LLaMA-3.1-8B 524288 327680 327680 524288 1179648 1179648 1179648 5242880 167772160
LLaMA-3.2-3B 393216 262144 262144 393216 720896 720896 720896 3473408 97255424
Mistral-7B-v0.3 524288 327680 327680 524288 1179648 1179648 1179648 5242880 167772160
@@ D.2 Commonsense Question-Answering Benchmark
We present the detailed accuracy scores on the commonsense question answering (CSQA) benchmark in this section, while the averaged scores are reported in Section 4. The results for each of the seven individual tasks, are provided in Table 11 and Table D.2, corresponding to Table 3.2 and Table LABEL:table:eval_detail, respectively.
@@
D.3 Ablation on Temperature $t$
We evaluate the effect of the temperature parameter $t$ in Equation 7 on the fine-tuned accuracy of LLaMA-3.2-3B under 4-bit quantization using the parameter budget $P(r=64)$. The results show that model accuracy remains stable for temperatures between 0.5 and 1.0, while performance slightly degrades when $t$ is too small or too large. A low temperature distributes the parameter budget nearly uniformly across channels, preventing sufficient allocation to channels with large quantization errors. Conversely, an overly high temperature over-concentrates parameters in these large-magnitude channels, neglecting important coefficients within low-magnitude channels and assigning unnecessary parameters within high-magnitude ones. As a result, both excessively low and excessively high temperatures lead to decreased fine-tuned model accuracy. Within the robust range of $t\in[0.5,1.0]$, we selected $t=1$ as the default setting, since it naturally favors allocating more parameters to outlier-included, large-magnitude channels while still maintaining stable empirical performance and methodological simplicity.
Table 13: Effect of temperature $t$ on GSM8k accuracy of LLaMA-3.2-3B under 4-bit quantization.
Temperature $t$ 0.25 0.5 1.0 1.5 2.0
GSM8k Acc. (%) 40.11 41.39 41.47 40.64 40.04
D.4 Ablation on Quantization Group Size
@@
Table 14: GSM8k accuracy (%) of QWHA on LLaMA-3.2-3B with 2-bit quantization using various quantization group sizes.
@@
D.5 Training Overhead of Adapters
Train time
@@ D.5 Training Overhead of Adapters
Memory Usage
@@
D.6 Initialization Overhead of Adapters
Initialization Time.
Table 17 reports the initialization latency of the low-rank adapter (CLoQ) and the FT-based adapters (QWHA and SSH) under the 4-bit setting across different models. The initialization of CLoQ requires gathering activations and quantization errors, followed by SVD decomposition before solving the least-squares problem. In contrast, the initialization of FT-based adapters involves collecting activations and quantization errors, followed by parameter selection and value assignment through channel-wise transforms and solving least-squares problem. During this process, the fast Hadamard kernel allows QWHA with WHT to perform efficient computation. As a result, QWHA achieves comparable initialization time to CLoQ, while SSH with the DHT kernel incurs significantly higher latency.
Table 17: Initialization latency (hours) of each method under the 4-bit setting.
Method LLaMA-3.2-3B LLaMA-3.1-8B Mistral-7B-v0.3
CLoQ 0.58 1.14 1.26
SSH 3.85 8.09 8.58
QWHA 0.66 1.34 1.46
Memory Usage.
We measure the peak memory consumption during initialization for each method on the LLaMA-3.2-3B model with 4-bit quantization, broken down into cached Hessians and model weights (Table 18). While every methods require multiple matrix projections, CLoQ additionally performs SVD, which leads to higher memory usage in layers with large dimensions. Overall, QWHA achieves slightly lower memory footprint due to its efficient block-wise computational implementation of the fast Hadamard transform.
Table 18: Memory usage (GB) and component breakdown during initialization of each method.
Method Total Usage Model Weight Cached Hessian
CLoQ 12.77 3.04 7.10
SSH 11.93 3.04 7.10
QWHA 11.52 3.04 7.10
D.7 Inference Overhead of Adapters
We investigate the inference throughput and memory usage of QWHA and CLoQ and present the results in Table 19 and Table 20, respectively. The evaluation uses a prefill length of 2048 and a generation length of 64, with batch size 128. We compare the FP16 pre-trained model, a quantized model with LoRA (corresponding to CLoQ), a sparse adapter (SHiRA), and FT-based adapters including WHA (QWHA), DCA (LoCA), and DHA (SSH).
As stated in SHiRA, sparse adapters provide a slight speedup over LoRA due to their simple scatter operations, compared to the low-rank matrix multiplications in LoRA. WHA introduces an additional WHT operation, but due to the fast Hadamard kernel, its overhead remains small: only a 1.9% decrease in throughput compared to LoRA. In contrast, conventional FT-based adapters such as DCA and DHA incur substantial overhead, showing a 50.9% throughput drop. Therefore, although QWHA applies an inverse WHT during inference, its overhead is marginal compared to LoRA, whereas other FT-based adapters experience significant efficiency degradation.
Regarding the inference memory usage, QWHA reduces overall memory usage by 13.0% compared to CLoQ, while both methods use 3.04 GB for model weights. This improvement arises from the use of a sparse adapter with efficient scatter operations. In addition, because the inverse WHT in QWHA is implemented without heavy matrix multiplications, it incurs no additional memory overhead and in fact results in lower peak memory usage. Consequently, QWHA achieves more than a 10% reduction in total memory consumption.
Table 19: Inference throughput (tokens/sec) of pretrained and quantized models of each adapters.
Method Pre-trained LoRA Sparse WHA DCA/DHA
Throughput (tokens/sec) 66.7 188.1 191.9 184.6 92.4
Table 20: Peak memory usage (GB) for each method.
Method QWHA CLoQ
Memory Usage (GB) 52.68 59.53
D.8 Training Curve
We analyzed the training loss, gradient norms, and convergence behavior throughout fine-tuning, and present the training curves and adapter gradient norms in Figure 10. We observe similar convergence behavior of QWHA and CLoQ. The gradient norms remain on a comparable scale despite QWHA’s large nominal scaling factor. This is because the effective scaling factor of both QWHA and CLoQ is close to 1.0. Moreover, under the same effective scaling value, low-rank adapters involve matrix multiplications during backpropagation, which naturally downscales the gradient norms applied to their parameters. In contrast, the sparse adapter in QWHA does not undergo this process, resulting in gradient norms that remain stable and consistently about twice as large throughout training.
Figure 10: Training loss (left) and gradient norm (right) of each methods during Alpaca fine-tuning in LLaMA-3.2-3B 4-bit model.